# Video-MME: The First-Ever Comprehensive Evaluation Benchmark of Multi-modal LLMs in Video Analysis

### v1

@@ Abstract
In the quest for artificial general intelligence, Multi-modal Large Language Models (MLLMs) have emerged as a focal point in recent advancements. However, the predominant focus remains on developing their capabilities in static image understanding. The potential of MLLMs in processing sequential visual data is still insufficiently explored, highlighting the absence of a comprehensive, high-quality assessment of their performance. In this paper, we introduce Video-MME, the first-ever full-spectrum, Multi-Modal Evaluation benchmark of MLLMs in Video analysis. Our work distinguishes from existing benchmarks through four key features: 1) Diversity in video types, spanning 6 primary visual domains with 30 subfields to ensure broad scenario generalizability; 2) Duration in temporal dimension, encompassing both short-, medium-, and long-term videos, ranging from 11 seconds to 1 hour, for robust contextual dynamics; 3) Breadth in data modalities, integrating multi-modal inputs besides video frames, including subtitles and audios, to unveil the all-round capabilities of MLLMs; 4) Quality in annotations, utilizing rigorous manual labeling by expert annotators to facilitate precise and reliable model assessment. 900 videos with a total of 256 hours are manually selected and annotated by repeatedly viewing all the video content, resulting in 2,700 question-answer pairs. With Video-MME, we extensively evaluate various state-of-the-art MLLMs, including GPT-4 series and Gemini 1.5 Pro, as well as open-source image models like InternVL-Chat-V1.5 and video models like LLaVA-NeXT-Video. Our experiments reveal that Gemini 1.5 Pro is the best-performing commercial model, significantly outperforming the open-source models with an average accuracy of 75.7%, compared to 52.5% for LLaVA-NeXT-Video. The results also demonstrate that Video-MME is a universal benchmark, which applies to both image and video MLLMs. Further analysis indicates that subtitle and audio information could significantly enhance video understanding. Besides, a decline in MLLM performance is observed as video duration increases for all models. Our dataset along with these findings underscores the need for further improvements in handling longer sequences and multi-modal data, shedding light on future MLLM development. Project page: https://video-mme.github.io.
@@ 1 Introduction
Using Video-MME, we benchmark various state-of-the-art MLLMs, including GPT-4V [45], GPT-4o [46], and Gemini 1.5 Pro [51], alongside open-source image models like InternVL-Chat-V1.5 [9] and video models like LLaVA-NeXT-Video [68]. Our experiments in Table 4 indicate that Gemini 1.5 Pro is the highest-performing commercial model, achieving an average accuracy of 75.7%. In comparison, open-source MLLMs exhibit substantial gaps compared to commercial models. For instance, the leading open-source model, LLaVA-NeXT-Video, attains an overall accuracy of 52.5%. These findings suggest there is considerable room for improvement in the open-source community. Our benchmark is also available to advanced image-based models by extending their input to multi-frame images, e.g., Qwen-VL-Max [5] and InternVL-Chat-V1.5 [9]. The accuracies of both the models reach 51% , which is close to that of the video specific model LLaVA-NeXT-Video, indicating that image understanding is the basis of video understanding, and the wide applicability of Video-MME in the filed of MLLMs. Further observations in Table 5 indicate that integrating subtitles and audios significantly enhances video comprehension capabilities, e.g., boosting Gemini 1.5 Pro by 5.9% and 4.7% respectively, with the gains being more pronounced for longer videos. A fine-grained analysis of task types reveals that subtitles and audios are particularly beneficial for videos requiring substantial domain knowledge. We also note a general decline in MLLM performance with increasing video length. This trend suggests that limitations in processing longer video sequences could be a critical bottleneck in the performance of MLLMs.
@@ 2.1 Dataset Construction
For a comprehensive coverage of different video types, we first create a domain hierarchy for collecting raw videos from YouTube. We define 6 key domains: Knowledge, Film & Television, Sports Competition, Life Record, and Multilingual, based on popular tendencies on YouTube. Each domain is further divided into detailed tags, such as football and basketball for sports competition, resulting in a total of 30 fine-grained video classes. The full domain-tag hierarchy and its distribution can be found in the left part of Figure 2. For each class, we collect videos with varying duration lengths, including short (less than 2 minutes), medium (4-15 minutes), and long videos (30-60 minutes). Besides, we also obtain corresponding meta-information such as subtitles (if provided) and audios for further investigation. Our final dataset consists of 900 videos spanning various domains with relatively balanced duration lengths, as depicted in the right part of Figure 2.
@@
After gathering the raw video data, we annotate it with high-quality question-answer (QA) pairs to evaluate the proficiency of MLLMs in interpreting video content. We employ a multiple-choice QA format to facilitate a straightforward and flexible assessment. Almost all authors of this study, proficient in English with extensive research experience in vision-language learning, perform the annotations. Specifically, they are first asked to watch the whole content of the video, and then to develop 3 corresponding questions, each with 4 potential options by repeatedly watching the video, contributing to 2,700 QA pairs in total. As shown in the bottom right corner of Figure 2, there are a total of 12 task types in the questions, including both perception, reasoning, and information synopsis. Particularly, each QA pair is required to be associated with the video content, avoiding MLLMs being able to answer without looking at the video.
@@ 2.2 Dataset Statistics
Our dataset comprises a total of 900 videos, 713 subtitles, and 869 audio files. Most videos are accompanied by both subtitles and audios, providing valuable resources for investigating the impact of external information on video understanding performance. The upper right part of Figure 2 illustrates the duration distribution of the collected videos. Specifically, within the short video category, longer videos occupy a larger portion. For medium-length videos, the duration is more uniformly distributed. In the category of long videos, there is a long-tailed distribution where longer videos have fewer samples. The bottom right part of Figure 2 shows the distribution of task types. Shorter videos predominantly involve perception-related tasks such as action and object recognition. In contrast, longer videos mainly feature tasks related to temporal reasoning. Overall, this analysis highlights that our dataset covers a wide range of video durations and various task types, enabling a comprehensive evaluation of temporal understanding.
@@
We demonstrate the language diversity of the questions and answers in our dataset. Table 2.2 lists the average word count of the textual fields in our dataset. The word counts for questions, options, and answers display notable consistency across different video lengths, suggesting a uniform style of QA pairs in our dataset. On the other hand, the word count for subtitles increases significantly with the length of the videos, e.g., short videos have an average word count of 200.9, while the long video subset have the count up to 6.5K. This trend indicates that longer videos contain more information, as evidenced by the increased volume of subtitles. The analysis reveals that our questions are diverse, and the answers are well-balanced. In addition, the distribution of the four answer options (A/B/C/D) follows a near-uniform distribution (25.2%/27.2%/25.2%/22.3%), ensuring an unbiased evaluation.
@@
Short 11.5 17.1 4.0 200.9
Medium 11.5 17.1 4.0 1450.7
Long 14.4 31.1 7.60 6515.3
@@
Inspired by EgoSchema [42], we adopt the certificate length to analyze the temporal difficulty of the QA pairs. The certificate of a given video QA pair is defined as the minimum set of sub-clips of the video that are both necessary and sufficient to convince a human verifier that the marked annotation is correct. The certificate length is calculated as the sum of the temporal lengths of the sub-clips identified. We randomly sample 5 videos from each class and calculate the certificate length distribution with extra annotators. As shown in Table 1, our dataset yields a median certificate length of 26.0s, 167.3s, and 890.7s for short, medium and long videos, respectively. Compared with the certificate length of EgoSchema, our medium and long video subset requires much longer video content digestion to answer the question. To the best of our knowledge, this analysis makes our Video-MME the most challenging Video QA dataset to date.
@@ 3 Experiments
Benchmarks #Videos #Clips Len.(s) #QA Pairs Anno. QA Tokens Sub. Tokens Multi-level Open-domain Sub.&Aud. MSRVTT-QA 2,990 2,990 15.2 72,821 A 8.4 MSVD-QA 504 504 9.8 13,157 A 7.6 TGIF-QA [18] 9,575 9,575 3.0 8,506 A&M 20.5 ActivityNet-QA [62] 800 800 111.4 8,000 M 10.2 TVQA [21] 2,179 15,253 11.2 15,253 M 27.8 159.8 How2QA 1,166 2,852 15.3 2,852 M 16.9 31.1 STAR 914 7,098 11.9 7,098 A 19.5 NExT-QA [57] 1,000 1,000 39.5 8,564 A 25.3 MVBench [24] 3,641 3,641 16.0 4,000 A 27.3 Video-Bench [44] 5,917 5,917 56.0 17,036 A&M 21.3 EgoSchema [42] 5,063 5,063 180.0 5,063 A&M 126.8 AutoEval-Video [8] 327 327 14.6 327 M 11.9 TempCompass [36] 410 500 11.4 7,540 A&M 49.2 Video-MME-S 300 300 80.8 900 M 28.6 200.9 Video-MME-M 300 300 520.2 900 32.8 1450.7 Video-MME-L 300 300 2471.0 900 45.4 6515.3 Video-MME 900 900 1024.0 2,700 35.6 3066.5
@@ 3.1 Settings
We conduct the evaluation on 3 commercial models, i.e., GPT-4V, GPT-4o [^1] Since the video interface of GPT-4o has not been released yet, we sample 10 frames and evaluate the model using multiple images as input., and Gemini 1.5 Pro, and 5 representative open-source video MLLMs, i.e., Video-LLaVA, VideoChat2, ST-LLM, Chat-UniVi-V1.5, and LLaVA-NeXT-Video. In addition, we also include advanced image MLLMs, i.e., Qwen-VL-Chat/Max and InternVL-Chat-V1.5, which usually have the ability to generalize to multi-image scenarios. We follow their official configurations and try to use more frames [^2] The numbers of the sampled frames are 10 for GPT-4V/o, 8 for Video-LLaVA, 16 for VideoChat2, 64 for ST-LLM, 32 for Chat-UniVi-V1.5, 32 for LLaVA-NeXT-Video, 4 for Qwen-VL-Chat/Max, and 10 for InternVL-Chat-V1.5. for evaluation. A special case is Gemini 1.5 Pro, because it supports extremely long multimodal contexts, so we take frames per second for both short and medium videos. For long videos, we capture a frame every two seconds to ensure the stability of the API testing. With respect to the setting of adding subtitles, all models except Gemini 1.5 Pro use the subtitles corresponding to the sampled video frames. Gemini 1.5 Pro uses all subtitles due to the full video frame sampling. In addition, only Gemini 1.5 Pro supports the input of audios by now, whose results are listed in Table 5.
@@ 3.2 Quantitative Results
Models LLM Params     Short (%) Medium (%) Long (%) Overall (%)         w/o subs w/ subs w/o subs w/ subs w/o subs w/ subs w/o subs w/ subs Open & Closed-source Image MLLMs Qwen-VL-Chat [5]     7B     46.4 47.1 38.1 39.8 38.0 38.3 40.9 41.7 Qwen-VL-Max [5]     -     56.5 58.3 49.9 49.8 49.0 46.9 51.8 51.7 InternVL-Chat-V1.5 [9]     20B     61.2 62.4 47.3 50.0 46.0 47.0 51.5 53.2 Open-source Video MLLMs Video-LLaVA [28]     7B     45.9 47.1 38.1 40.2 37.3 39.6 40.4 42.3 VideoChat2 [24]     7B     38.2 41.6 33.2 34.3 29.7 31.9 33.7 35.9 ST-LLM [33]     7B     47.0 49.9 36.9 42.2 31.8 37.3 38.6 43.2 Chat-UniVi-V1.5 [19]     7B     46.3 51.4 40.3 45.2 36.9 42.3 41.2 46.3 LLaVA-NeXT-Video [68]     34B     63.1 66.4 51.1 53.2 44.6 48.7 52.5 56.0 Closed-source MLLMs GPT-4V [45]     -     71.4 74.5 56.5 59.3 54.2 57.2 60.7 63.7 GPT-4o [46]     -     77.1 77.5 62.1 63.0 59.2 56.7 66.2 65.8 Gemini 1.5 Pro [51]     -     82.3 84.7 75.3 82.6 67.5 76.3 75.7 81.6
Table 5: Performance of Gemini 1.5 Pro on six major categories of Video-MME. The input modality includes frames only, frames with subtitles, and frames with audios.
Subset Modality Category Knowledge Film & Television Sports Artistic Life Multilingual Overall Competition Performance Record Short Frames 79.8 82.5 78.7 88.6 86.7 70.0 82.3 + Subs 85.6 (+5.8) 87.5 (+5.0) 77.3 (-1.4) 87.6 (-1.0) 85.7 (-1.0) 86.7 (+16.7) 84.7 (+2.4) + Audios 84.8 (+5.0) 85.0 (+2.5) 78.7 (+0.0) 86.7 (-1.9) 87.1 (+0.4) 83.3 (+13.3) 84.5 (+2.2) Medium Frames 70.0 82.8 73.9 84.3 71.1 95.8 75.3 + Subs 84.3 (+14.3) 89.9 (+7.1) 78.3 (+4.4) 84.3 (-) 77.5 (+6.4) 91.7 (+4.1) 82.6 (+7.3) + Audios 80.5 (+10.5) 87.5 (+4.7) 73.9 (-) 82.4(-1.9) 78.5 (+7.4) 100.0 (+4.2) 80.6 (+5.3) Long Frames 74.6 70.1 58.1 60.0 66.7 66.7 67.5 + Subs 82.1 (+7.5) 73.6 (+2.5) 72.6 (+14.5) 70.0 (+10.0) 75.6 (+8.9) 75.0 (+8.3) 76.3 (+8.8) + Audios 81.9 (+7.3) 75.6 (+5.5) 71.9 (+13.8) 66.7 (+6.7) 67.2 (+0.5) 83.3 (+16.6) 74.7 (+7.2) Overall Frames 75.0 79.1 71.1 78.5 76.5 76.9 75.7 +Subs 84.0 (+9.0) 84.3 (+5.2) 76.3 (+5.2) 81.1 (+2.6) 80.4 (+3.9) 84.6 (+7.7) 81.6 (+5.9) +Audios 82.5 (+7.5) 83.2 (+4.1) 75.1 (+4.0) 79.1 (+0.6) 79.4 (+2.9) 88.5 (+11.6) 80.4 (+4.7)
@@
As one of the pioneering commercial large models integrated with video comprehension capabilities, Gemini 1.5 Pro has achieved the best performance among its peers on Video-MME. As depicted in Figure 3, with video frames as input alone, Gemini 1.5 Pro attains an accuracy of 75.7%, surpassing GPT-4V and GPT-4o by 15% and 9.5%, respectively. Table 5 shows the fine-grained performance of Gemini 1.5 Pro. Among the six major video categories, Gemini 1.5 Pro performs the best in Film & Television, achieving an accuracy of 79.1%, while performing the lowest in the Sports Competition category, with an accuracy of 71.1%. As video duration increases, Gemini 1.5 Pro’s performance declines (e.g., --14.8% from short to long videos), highlighting the model’s weakness in capturing long-range temporal relationships. Nevertheless, Gemini 1.5 Pro’s performance on long videos still surpasses all open-source models on short videos, demonstrating its superior capabilities. In addition to visual frame input, Gemini 1.5 Pro’s support for additional modalities, including subtitles and audios, provides opportunities for further performance improvement. For example, Table 5 displays that using audios can increase accuracy by 7.2% for long videos, the improvement in the multilingual category even reaches 16.6%. We can also see that the effect of subtitles and audios is different in these six categories. These motivate future research to develop versatile models that can support a wider range of modality inputs.
@@
As shown in Table 4, among the 7B models, Chat-UniVi-V1.5 achieves the best performance with 41.2%. LLaVA-NeXT-Video with 34B LLM achieves an accuracy of 52.5%, demonstrating its stronger capabilities, especially in the tasks of attribute perception, spatial reasoning, and information synopsis, as exhibited in Figure 3. Nevertheless, there still remains a significant gap between LLaVA-NeXT-Video and Gemini Pro 1.5, particularly in counting problems, action recognition, and temporal perception, indicating substantial room for improvement. It is observed that adding subtitles can also help the open-sourced models. For example, the accuracy of LLaVA-NeXT-Video improves from 52.5% to 56.0%. It is regrettable that none of the open-sourced models support audio input.

### v2

@@ Abstract
In the quest for artificial general intelligence, Multi-modal Large Language Models (MLLMs) have emerged as a focal point in recent advancements. However, the predominant focus remains on developing their capabilities in static image understanding. The potential of MLLMs in processing sequential visual data is still insufficiently explored, highlighting the absence of a comprehensive, high-quality assessment of their performance. In this paper, we introduce Video-MME, the first-ever full-spectrum, Multi-Modal Evaluation benchmark of MLLMs in Video analysis. Our work distinguishes from existing benchmarks through four key features: 1) Diversity in video types, spanning 6 primary visual domains with 30 subfields to ensure broad scenario generalizability; 2) Duration in temporal dimension, encompassing both short-, medium-, and long-term videos, ranging from 11 seconds to 1 hour, for robust contextual dynamics; 3) Breadth in data modalities, integrating multi-modal inputs besides video frames, including subtitles and audios, to unveil the all-round capabilities of MLLMs; 4) Quality in annotations, utilizing rigorous manual labeling by expert annotators to facilitate precise and reliable model assessment. 900 videos with a total of 254 hours are manually selected and annotated by repeatedly viewing all the video content, resulting in 2,700 question-answer pairs. With Video-MME, we extensively evaluate various state-of-the-art MLLMs, including GPT-4 series and Gemini 1.5 Pro, as well as open-source image models like InternVL-Chat-V1.5 and video models like LLaVA-NeXT-Video. Our experiments reveal that Gemini 1.5 Pro is the best-performing commercial model, significantly outperforming the open-source models with an average accuracy of 75%, compared to 71.9% for GPT-4o. The results also demonstrate that Video-MME is a universal benchmark, which applies to both image and video MLLMs. Further analysis indicates that subtitle and audio information could significantly enhance video understanding. Besides, a decline in MLLM performance is observed as video duration increases for all models. Our dataset along with these findings underscores the need for further improvements in handling longer sequences and multi-modal data, shedding light on future MLLM development. Project page: https://video-mme.github.io.
@@ 1 Introduction
Using Video-MME, we benchmark various state-of-the-art MLLMs, including GPT-4V [48], GPT-4o [49], and Gemini 1.5 Pro [54], alongside open-source image models like InternVL-Chat-V1.5 [9] and video models like LLaVA-NeXT-Video [74]. Our experiments in Table 4 indicate that Gemini 1.5 Pro is the highest-performing commercial model, achieving an average accuracy of 75%. In comparison, open-source MLLMs exhibit substantial gaps compared to commercial models. For instance, the leading open-source model, VILA-1.5 [31], attains an overall accuracy of 59%. These findings suggest there is considerable room for improvement in the open-source community. Our benchmark is also available to advanced image-based models by extending their input to multi-frame images, e.g., Qwen-VL-Max [5] and InternVL-Chat-V1.5 [9]. The accuracies of both the models reach 50% , which is close to that of the video specific model LLaVA-NeXT-Video, indicating that image understanding is the basis of video understanding, and the wide applicability of Video-MME in the filed of MLLMs. Further observations in Table 5 indicate that integrating subtitles and audios significantly enhances video comprehension capabilities, e.g., boosting Gemini 1.5 Pro by 6.2% and 4.3% respectively, with the gains being more pronounced for longer videos. A fine-grained analysis of task types reveals that subtitles and audios are particularly beneficial for videos requiring substantial domain knowledge. We also note a general decline in MLLM performance with increasing video length. This trend suggests that limitations in processing longer video sequences could be a critical bottleneck in the performance of MLLMs.
@@ 2.1 Dataset Construction
For a comprehensive coverage of different video types, we first create a domain hierarchy for collecting raw videos from YouTube. We define 6 key domains: Knowledge, Film & Television, Sports Competition, Life Record, and Multilingual, based on popular tendencies on YouTube. Each domain is further divided into detailed tags, such as football and basketball for sports competition, resulting in a total of 30 fine-grained video classes. The full domain-tag hierarchy and its distribution can be found in the left part of Figure 2. For each class, we collect videos with varying duration lengths, including short (< 2 minutes), medium (4-15 minutes), and long videos (30-60 minutes). Besides, we also obtain corresponding meta-information such as subtitles (if provided) and audios for further investigation. Our final dataset consists of 900 videos with 744 subtitles and 900 audios, spanning various domains with relatively balanced duration lengths, as depicted in the right part of Figure 2.
@@
After gathering the raw video data, we annotate it with high-quality question-answer (QA) pairs to evaluate the proficiency of MLLMs in interpreting video content. We employ a multiple-choice QA format to facilitate a straightforward and flexible assessment. The researchers proficient in English with extensive research experience in vision-language learning, perform the annotations. Specifically, they are first asked to watch the whole content of the video, and then to develop 3 corresponding questions, each with 4 potential options by repeatedly watching the video, contributing to 2,700 QA pairs in total. As shown in the bottom right corner of Figure 2, there are a total of 12 task types in the questions, including both perception, reasoning, and information synopsis. Particularly, each QA pair is required to be associated with the video content, avoiding MLLMs being able to answer without looking at the video.
@@ 2.2 Dataset Statistics
Our dataset comprises a total of 900 videos, 744 subtitles, and 900 audio files. Most videos are accompanied by both subtitles and audios, providing valuable resources for investigating the impact of external information on video understanding performance. The upper right part of Figure 2 illustrates the duration distribution of the collected videos. Specifically, within the short video category, longer videos occupy a larger portion. For medium-length videos, the duration is more uniformly distributed. In the category of long videos, there is a long-tailed distribution where longer videos have fewer samples. The bottom right part of Figure 2 shows the distribution of task types. Shorter videos predominantly involve perception-related tasks such as action and object recognition. In contrast, longer videos mainly feature tasks related to temporal reasoning. Overall, this analysis highlights that our dataset covers a wide range of video durations and various task types, enabling a comprehensive evaluation of temporal understanding.
@@
We demonstrate the language diversity of the questions and answers in our dataset. Table 2.2 lists the average word count of the textual fields in our dataset. The word counts for questions, options, and answers display notable consistency across different video lengths, suggesting a uniform style of QA pairs in our dataset. On the other hand, the word count for subtitles increases significantly with the length of the videos, e.g., short videos have an average word count of 198.6, while the long video subset have the count up to 6.5K. This trend indicates that longer videos contain more information, as evidenced by the increased volume of subtitles. The analysis reveals that our questions are diverse, and the answers are well-balanced. In addition, the distribution of the four answer options (A/B/C/D) follows a near-uniform distribution (25.1%/27.2%/25.3%/22.4%), ensuring an unbiased evaluation.
@@
Short 11.5 17.2 4.0 198.6
Medium 12.2 20.6 5.0 1425.6
Long 14.5 31.0 7.5 6515.6
@@
Inspired by EgoSchema [45], we adopt the certificate length to analyze the temporal difficulty of the QA pairs. The certificate of a given video QA pair is defined as the minimum set of sub-clips of the video that are both necessary and sufficient to convince a human verifier that the marked annotation is correct. The certificate length is calculated as the sum of the temporal lengths of the sub-clips identified. We randomly sample 3 videos from each class and calculate the certificate length distribution with extra annotators. As shown in Table 1, our dataset yields a median certificate length of 26s, 164.7s, and 890.7s for short, medium and long videos, respectively. Compared with the certificate length of EgoSchema, our medium and long video subset requires much longer video content digestion to answer the question. To the best of our knowledge, this analysis makes our Video-MME the most challenging Video QA dataset to date.
@@ 3 Experiments
Benchmarks #Videos #Clips Len.(s) #QA Pairs Anno. QA Tokens Sub. Tokens Multi-level Open-domain Sub.&Aud. MSRVTT-QA [63] 2,990 2,990 15.2 72,821 A 8.4 MSVD-QA [63] 504 504 9.8 13,157 A 7.6 TGIF-QA [18] 9,575 9,575 3.0 8,506 A&M 20.5 ActivityNet-QA [68] 800 800 111.4 8,000 M 10.2 TVQA [21] 2,179 15,253 11.2 15,253 M 27.8 159.8 How2QA [25] 1,166 2,852 15.3 2,852 M 16.9 31.1 STAR [61] 914 7,098 11.9 7,098 A 19.5 NExT-QA [62] 1,000 1,000 39.5 8,564 A 25.3 MVBench [24] 3,641 3,641 16.0 4,000 A 27.3 Video-Bench [47] 5,917 5,917 56.0 17,036 A&M 21.3 EgoSchema [45] 5,063 5,063 180.0 5,063 A&M 126.8 AutoEval-Video [8] 327 327 14.6 327 M 11.9 TempCompass [39] 410 500 11.4 7,540 A&M 49.2 Video-MME-S 300 300 80.7 900 M 28.7 198.6 Video-MME-M 300 300 515.9 900 32.8 1425.6 Video-MME-L 300 300 2466.7 900 45.6 6515.6 Video-MME 900 900 1017.9 2,700 35.7 3086.5
@@ 3.1 Settings
We conduct the evaluation on 4 commercial models, i.e., GPT-4V, GPT-4o, Gemini 1.5 Flash, and Gemini 1.5 Pro. Representative open-source video MLLMs including Video-LLaVA, VideoChat2-Mistral, ST-LLM, Chat-UniVi-V1.5, LLaVA-NeXT-Video, and VILA-1.5 are evaluated as well. In addition, we also include advanced image MLLMs, i.e., Qwen-VL-Chat/Max and InternVL-Chat-V1.5, which usually can generalize to multi-image scenarios. We follow their official configurations and try to use more frames [^1] The numbers of the sampled frames are 10 for GPT-4V, 384 for GPT-4o, 8 for Video-LLaVA, 16 for VideoChat2-Mistral, 64 for ST-LLM, 64 for Chat-UniVi-V1.5, 32 for LLaVA-NeXT-Video, 8 for VILA-1.5, 4 for Qwen-VL-Chat/Max, and 10 for InternVL-Chat-V1.5. for evaluation. A special case is Gemini 1.5 Pro, because it supports extremely long multimodal contexts, so we take frames per second for both short and medium videos. For long videos, we capture a frame every two seconds to ensure the stability of the API. With respect to the setting of adding subtitles, all models except Gemini 1.5 Pro use the subtitles corresponding to the sampled video frames. For example, if you sample 10 frames, take the 10 subtitles that correspond to the time of those 10 frames. Gemini 1.5 Pro uses all subtitles due to the full video frame sampling. Besides, only Gemini 1.5 Pro supports the input of audios by now, whose results are listed in Table 5.
@@ 3.2 Quantitative Results
Models LLM Params     Short (%) Medium (%) Long (%) Overall (%)         w/o subs w/ subs w/o subs w/ subs w/o subs w/ subs w/o subs w/ subs Open & Closed-source Image MLLMs Qwen-VL-Chat [5]     7B     46.9 47.3 38.7 40.4 37.8 37.9 41.1 41.9 Qwen-VL-Max [5]     -     55.8 57.6 49.2 48.9 48.9 47.0 51.3 51.2 InternVL-Chat-V1.5 [9]     20B     60.2 61.7 46.4 49.1 45.6 46.6 50.7 52.4 Open-source Video MLLMs Video-LLaVA [30]     7B     45.3 46.1 38.0 40.7 36.2 38.1 39.9 41.6 ST-LLM [36]     7B     45.7 48.4 36.8 41.4 31.3 36.9 37.9 42.3 VideoChat2-Mistral [24]     7B     48.3 52.8 37.0 39.4 33.2 39.2 39.5 43.8 Chat-UniVi-V1.5 [19]     7B     45.7 51.2 40.3 44.6 35.8 41.8 40.6 45.9 LLaVA-NeXT-Video [74]     34B     61.7 65.1 50.1 52.2 44.3 47.2 52.0 54.9 VILA-1.5 [31]     34B     68.1 68.9 58.1 57.4 50.8 52.0 59.0 59.4 Closed-source MLLMs GPT-4V [48]     -     70.5 73.2 55.8 59.7 53.5 56.9 59.9 63.3 GPT-4o [49]     -     80.0 82.8 70.3 76.6 65.3 72.1 71.9 77.2 Gemini 1.5 Flash [54]     -     78.8 79.8 68.8 74.7 61.1 68.8 70.3 75.0 Gemini 1.5 Pro [54]     -     81.7 84.5 74.3 81.0 67.4 77.4 75.0 81.3
Table 5: Performance of Gemini 1.5 Pro on 6 major categories of Video-MME. The input modality includes frames only, frames with subtitles, and frames with audios.
Subset Modality Category Knowledge Film & Television Sports Artistic Life Multilingual Overall Competition Performance Record Short Frames 78.3 80.8 76.7 86.7 88.1 76.7 81.7 + Subs 83.3 (+4.9) 86.7 (+5.8) 79.3 (+2.7) 87.6 (+1.0) 86.6 (-1.5) 86.7 (+10.0) 84.5 (+2.8) + Audio 81.4 (+3.1) 87.5 (+6.7) 78.7 (+2.0) 86.7 (-) 85.6 (-2.4) 86.7 (+10.0) 83.6 (+1.9) Medium Frames 70.2 81.2 68.7 84.3 73.4 87.5 74.3 + Subs 83.3 (+13.1) 84.9 (+3.7) 76.2 (+7.5) 85.3 (+1.0) 76.8 (+3.4) 83.3 (-4.2) 81.0 (+6.7) + Audio 80.2 (+9.9) 83.9 (+2.6) 72.1 (+3.4) 84.3 (-) 76.8 (+3.4) 100.0 (+12.5) 79.5 (+5.2) Long Frames 73.4 70.1 58.3 63.3 65.1 70.8 67.4 + Subs 83.0 (+9.6) 71.4 (+1.3) 77.5 (+19.2) 70.0 (+6.7) 74.6 (+9.5) 87.5 (+16.7) 77.4 (+10.1) + Audio 81.1 (+7.7) 73.2 (+3.1) 72.6 (+14.3) 63.3 (-) 66.7 (+1.6) 83.3 (+12.5) 73.6 (+6.2) Overall Frames 74.1 77.9 68.6 78.8 77.4 78.2 75.0 + Subs 83.2 (+9.2) 81.8 (+3.9) 77.7 (+9.1) 81.5 (+2.7) 80.3 (+2.9) 85.9 (+7.7) 81.3 (+6.2) + Audio 80.9 (+6.8) 82.4 (+4.5) 74.6 (+6.1) 78.8 (-) 78.0 (+0.6) 89.7 (+11.5) 79.4 (+4.3)
@@
As one of the pioneering commercial large models integrated with video comprehension capabilities, Gemini 1.5 Pro has achieved the best performance among its peers on Video-MME. As depicted in Table 4, with video frames as input alone, Gemini 1.5 Pro attains an accuracy of 75%, surpassing GPT-4V and GPT-4o by 15.1% and 3.1%, respectively. Table 5 shows the fine-grained performance of Gemini 1.5 Pro. Among the 6 major video categories, Gemini 1.5 Pro performs the best in Artistic Performance, achieving an accuracy of 78.8%, while performing the lowest in the Sports Competition category, with an accuracy of 68.6%. As video duration increases, Gemini 1.5 Pro’s performance declines (e.g., --6.7% from short to long videos), highlighting the model’s weakness in capturing long-range temporal relationships. Nevertheless, Gemini 1.5 Pro’s performance on long videos still surpasses almost all open-source models, except for VILA-1.5, on short videos, demonstrating its superior capabilities. In addition to visual frame input, Gemini 1.5 Pro’s support for additional modalities, including subtitles and audios, provides opportunities for further performance improvement. For example, Table 5 displays that using audios can increase accuracy by 6.2% for long videos, and the improvement in the multilingual category even reaches 16.7%. We can also see that the effect of subtitles and audios is different in these six categories. These motivate future research to develop versatile models that can support a wider range of modality inputs.
@@
As shown in Table 4, among the 7B models, Chat-UniVi-V1.5 achieves the best performance with 40.6%. VILA-1.5 with 34B LLM achieves an accuracy of 59%, demonstrating its stronger capabilities, especially in the tasks of spatial reasoning, attribute perception, and information synopsis, as exhibited in Figure 3. Nevertheless, there still remains a significant gap between VILA-1.5 and Gemini Pro 1.5, particularly in counting problems, action recognition, and temporal perception, indicating substantial room for improvement. It is observed that adding subtitles can also help the open-sourced models. For example, the accuracy of LLaVA-NeXT-Video improves from 52% to 54.9%. It is regrettable that none of the open-sourced models support audio input.